Five Seasons Hotel replenishes 30,000 bars of high-quality soaps per year for its hotel rooms.

The ordering cost is $50 per order.

The annual carrying cost is $8.25 per unit.

The lead time is 5 days.

The hotel operates 365 days a year.

In order to protect the hotel against uncertainty in demand for soaps, the hotel maintains a safety stock of 300 bars of high-quality soaps.

a. What is the Total Annual Inventory Cost?

b. What is the Total Number of Orders placed per Year?

c. What is the Re-Order Cycle?

d. What must be the hotel’s minimum inventory level for those high-quality soaps?

e. What should be the hotel’s maximum inventory level for those high-quality soaps?
Answer #1

Answer:

Days per year= 365
Annual demand, D= 30000
Order cost or setup cost, S $50.0
holding cost per year, H= $8.250
Lead time, LT 5.00
Economic Order quantity, Q= squareroot(2*S*D/H) 603.0
Answer b #orders per year = D/Q 49.75
annual holding cost= Q*H/2 $2,487.47
Annual ordering cost= D*S/Q $2,487.47
Answer a total ordering+holding cost $4,974.94
Answer c reorder cycle= days per year/orders per year 7.34 days

Answer d: min. inventory level is the safety stock of 300 units

Answer e: Max inventory level, is when the order arrives= 300+ 603= 903